OpenStudy (anonymous):

I don't understand factoring, so I'm so confused on this. Find the zeros of the function f(x) = x^2 - 5x - 6 by factoring.

hartnn (hartnn):

\(Basically, \\ \large ☻☺+☻♥= ☻(☺+♥) \\ \text{this is factoring, } \\ \text{☻ was common from both terms and factored out.}\)

OpenStudy (anonymous):

It shows nothing but question marks? @hartnn

hartnn (hartnn):

refrsh your page :)

OpenStudy (anonymous):

Oh! Ahh fixed I see it!! So the two xs would be out and left with the negative six?

hartnn (hartnn):

factoring trinomials doesn't exactly work that way. We split the middle term first and then factor out whatever we can from the 2 pairs of terms we get after splitting... let me give you an example :)

OpenStudy (mrnood):

te reason why we use factoring to fin zeros is because if you multiply 2 numbers together and the answer is zero, then at least ONE of those numbers must be zero That is if a x b = 0 then either a=0, OR b=0 (or a and b are BOTH zero) SO If you star with the equation and want to find its zeros then it would be good if we can write the equation as (M) x (N) Note that not ALL quadratics can be factorised. so - the equation has x^2 in it so it must be made up of (x) x (x) the equation also has a constant term without any x in it which must be made up of multiplying two constant terms so we start with the assumption that your equation can be re written as (x+a)(x+b) = 0 If oyu expand those brackets you get x^2 + x(a+b) + ab = 0 and you can see that comparing this to your equation a+b = -5 ab= 6 so you need to work out a & b so those two things are true. so if ab = 6 then a & b could be: 1,6 6,1 -1, -6 -6, -1 3,2 2,3 -2,-3 -3-2 these are the only options so now look at those and find which pair ADD up to -5

OpenStudy (anonymous):

Thats a lot of possible answers @MrNood

hartnn (hartnn):

\(x^2+6x+8\) We find 2 numbers whose sum is +6 and product is +8 Such 2 numbers are easy to find, here, +4 and +2 so we split the middle term like +6x= +4x+2x Thats the first step of splitting...

OpenStudy (mrnood):

Those are the [possible ways of getting 6 from multiplying two numbers Just take a SIMPLE look at them, and firn WHICH pair adds up to -5 There is only 1

OpenStudy (anonymous):

But couldn't it be others as well? How do you find the right one?? @hartnn

OpenStudy (mrnood):

as I said - those are the ONLY ways of multiplying two numbers together to make 6, and ONLY one of them adds up to -5 Take a look, and find which pair adds up to -5

OpenStudy (anonymous):

-3,-2 @MrNood

OpenStudy (mrnood):

OOPs - sorry - the number we want is -6, not 6 the options are: -1,6 -6,1 -3,2 -2,3 sorry - same issue find the ones that add to -5

OpenStudy (anonymous):

Okay I semi get it. I'll try the next one with this system. Thanks @MrNood & @hartnn